# Daejeon 34141, South Korea

### v1

@@ Abstract.
Let $N$ be a manifold of dimension $m$ with a flat vector bundle given by a representation $\rho:\pi_{1}(N)\rightarrow{\mathsf{GL}}(n,{\mathbb{R}})$ where $\pi_{1}(N)$ is finitely generated. The holonomy group $\rho$ is a $k$-partially hyperbolic holonomy representation if the flat bundle pulled back over the unit tangent bundle of a sufficiently large compact submanifold of $N$ splits into expanding, neutral, and contracting subbundles along the geodesic flow, where the expanding and contracting subbundles are $k$-dimensional with $k<n/2$. Suppose that each element of $\rho(\pi_{1}(N))$ has an eigenvalue of norm $1$, or, alternatively, each element of it has a singular value that is uniformly bounded above and below. We show that $\rho$ is a $P$-Anosov representation for a parabolic subgroup $P$ of ${\mathsf{GL}}(n,{\mathbb{R}})$ if and only if $\rho$ is a partially hyperbolic representation. We are going to primarily employ representation theory techniques. As an application, we will show this holds when $N$ is a complete affine $n$-manifold and $\rho$ is a linear part of the holonomy representation. This had never been done over the full general linear group.
@@ 1.1. Notation
Let $N$ be an $m$-manifold with a finitely generated fundamental group $\pi_{1}(N)$. Each flat vector bundle over $N$ can be described as ${\mathbb{R}}^{n}_{\rho}(N):=\tilde{N}\times{\mathbb{R}}^{n}/\sim$ where $(x,\vec{v})\sim(\gamma(x),\rho(\gamma)(\vec{v})$ for $\gamma\in\pi_{1}(N)$ where $\rho:\pi_{1}(N)\rightarrow{\mathsf{GL}}(n,{\mathbb{R}})$ is a representation. A differential-geometry description of a flat bundle is a vector bundle with a flat linear connection.
Let $M$ be a compact submanifold (with or without boundary) in $N$. We assume that the inclusion map $M\hookrightarrow N$ induces a surjection on fundamental groups. The inverse image $\hat{M}$ of $M$ in $\tilde{N}$ is connected, and the deck transformation group $\Gamma_{M}$ of $\hat{M}$ is isomorphic to $\Gamma$. In these cases, we refer to $M$, the map $\hookrightarrow N, M^\hat{M}$, the projection $p_{\hat{M}}:\hat{M}\to M$, and $\Gamma_{M}$ as a fundamental group surjective (FS) submanifold, an FS map, an FS cover, an FS covering map, and an FS deck transformation group, respectively.
@@
We will assume that $\partial M$ is convex and smooth in $N$. Let $d_{M}$ denote the path metric induced from a Riemannian metric on $N$, and let $d_{\hat{M}}$ denote the path metric on $\hat{M}$ induced from it.
  • Let $\mathbf{U}M$ denote the space of direction vectors on $M$, i.e., the space of elements of the form $(x,\vec{v})$ for $x\in M$ and a unit vector $\vec{v}\in T_{x}M-\{O\}$ with the projection $\pi_{\mathbf{U}M}:\mathbf{U}M\rightarrow M$.
@@
A complete isometric geodesic in $\hat{M}$ is a geodesic that is an isometry of $\mathbb{R}$} into $\hat{M}$ equipped with $M$. Some but not all Riemannian geodesics are isometries. Actually, in the geodesic metric spaces, these are identical to the notion of geodesics. Since we are working on Riemannian metric spaces, we use the adjectives. A complete isometric geodesic in $M$ is a geodesic that lifts to a complete isometric geodesic in $\hat{M}$. A isometric ray in $\hat{M}$ is an isometric geodesic defined on $0,\infty)$.
@@
$\gamma\cdot(x,\vec{v})=(\gamma(x),\rho(\gamma)(\vec{v}))\hbox{ for }x\in\mathbf{UC}\hat{M},\vec{v}\in{\mathbb{R}}^{n}.$
@@ 1.2. Main result
In this paper, we establish connections between the partial hyperbolicity in the context of dynamical systems and the P-Anosov properties.
@@
    ${\mathbb{V}}_{+}(x),{\mathbb{V}}_{0}(x),{\mathbb{V}}_{-}(x)\subset{\mathbb{V}}(x)$
    for the fiber ${\mathbb{V}}(x)$ of $\mathbb{V}$} over each point $x$ of $\mathbf{UC}M$ invariant under the flow $\Phi_{t}$ of $\mathbb{V}$} over the geodesic flow $\phi_{t}$ of $\mathbf{UC}M$.
@@
    ${\mathbb{V}}_{+}(x),{\mathbb{V}}_{0}(x)$ and ${\mathbb{V}}_{-}(x)$ for each $x\in\mathbf{UC}M$ are independent subspaces, and their sum equals ${\mathbb{V}}(x)$.
@@
We additionally require that $\dim{\mathbb{V}}_{+}=\dim{\mathbb{V}}_{-}\geq 1$. Here $\dim{\mathbb{V}}_{+}$ is the partial hyperbolicity index of $\rho$ for the associated bundle decomposition ${\mathbb{V}}_{+},{\mathbb{V}}_{0},{\mathbb{V}}_{-}$ where $\dim{\mathbb{V}}_{+}$ is maximal and $\dim{\mathbb{V}}_{+}=\dim{\mathbb{V}}_{-}<n/2$, so that $\dim{\mathbb{V}}_{0}\geq 1$. Furthermore, ${\mathbb{V}}_{0}$ is said to be the neutral subbundle of $\mathbb{V}$}. (See Definition 1.5 of [10] or [6]. We use the words “in a bundle sense” only as a convenience of this paper.)
The definition depends on the Riemannian metric, but not on $\left|\mkern-2.0mu\left|\cdot\right|\mkern-2.0mu\right|_{{\mathbb{R}}^{n}_{\rho}}$. However, if $\Gamma_{M}$ is word hyperbolic, then we can replace $\mathbf{UC}\hat{M}$ with a Gromov flow space which depends only on $\Gamma_{M}$ up to quasi-isometry. (See Theorem 2.6 and Gromov [23] or Mineyev [36].)
@@
  • the linear part $\rho:\pi_{1}(N)=\Gamma_{M}\to{\mathsf{GL}}(n,\mathbb{R})$ is a partially hyperbolic representation in a bundle sense over $M$.
Since partial hyperbolicity is stronger than the Anosov property by Theorem 1.1, $\pi_{1}(N)$ is always hyperbolic. By Theorem 2.6, there exists a quasi-isometry from $\mathbf{UC}\hat{M}$ to the Gromov flow space $\partial^{(2)}_{\infty}M\times\mathbb{R}$ preserving the geodesic length parameter. Hence, the discussion reduces to one on the Gromov flow space, which is independent of the choice of FS submanifolds.
We denote by $a_{1}(A),\ldots,a_{n}(A)$ the nonincreading set of singular values of an $n\times n$-matrix $A$. We denote by $\lambda_{1}(A),\ldots,\lambda_{n}(A)$ the nonincreasing set of norms of eigenvalues of an $n\times n$-matrix $A$. (Here, the multiplicity is allowed.)
@@
$\theta=\left\{\log\lambda_{i_{1}}-\log\lambda_{i_{1}+1},\ldots,\log\lambda_{i_{m}}-\log\lambda_{i_{m}+1}\right\},1\leq i_{1}<\cdots<i_{m}\leq n-1,$
of $\mathrm{GL}(n,{\mathbb{R}})$, the parabolic subgroup $P_{\theta}$ (resp. $\left.P_{\theta}^{-}\right)$ is the subgroup of block upper (resp. lower) triangular matrices in $\mathrm{GL}(n,{\mathbb{R}})$ with square diagonal blocks of sizes $i_{1},i_{2}-i_{1},\ldots,i_{m}-i_{m-1},n-i_{m}$ respectively.
@@
Let $N$ be an $m$-manifold with a finitely generated fundamental group. Let $\rho:\pi_{1}(N)\rightarrow{\mathsf{GL}}(n,{\mathbb{R}})$ be a representation. Assume one of the following:
  • $|\log\lambda_{i_{g}}(\rho(g))|=1$ for all $g\in\pi_{1}(M)$ where the index $i_{g}\in\{1,\dots,n\}$ may depend on $g$.
  • There exists a constant $C>0$ such that $|\log a_{j_{g}}(\rho(g))|<C$ for all $g\in\pi_{1}(M)$ where the index $j_{g}\in\{1,\dots,n\}$ may depend on $g$.
Then $\rho$ is $P$-Anosov for any parabolic subgroup $P$ of ${\mathsf{GL}}(n,{\mathbb{R}})$ with $P=P_{\theta}$ for $\theta$ containing $\log\lambda_{k}-\log\lambda_{k+1},k\leq n/2$ if and only if $\rho$ is partially hyperbolic of the index $k$ in a bundle sense. In these cases $k<n/2$.
@@
We can ask if an Anosov representation is always partially hyperbolic up to multiplying by a positive scalar representation to $\mathbb{R}$}. We cannot answer this question since controlling the singular values of representation elements is very hard in general, and the only related results that we know of are due to Abel-Margulis-Soifer [1], which is not applicable here as far as we know.
@@
Suppose that $N$ is a complete affine $n$-manifold with a finitely generated fundamental group. Let $\rho:\pi_{1}(N)\rightarrow{\mathsf{GL}}(n,{\mathbb{R}})$ be a linear part of the affine holonomy group. Then the conclusions of Theorem 1.1 hold.
@@
We wish to approach the conjecture that a closed complete affine manifold cannot have a word hyperbolic fundamental group. One approach considered is using Osedelec-type theorems (See Chapter 8 of [31]). However, generalizing our results to measurable settings is currently hindered by numerous technical issues. Discussions on these matters took place with researchers such as Nicholas Tholozan at BIRS, Banff, during the winter of 2020.
@@ 1.3. Some historical comments on complete affine manifolds
Further developments include Ghosh [15, 17, 16, 18], Ghosh-Treib [19], and Smilga [38, 39, 40, 41, 42], who generalized these results and produced higher-dimensional examples of proper and free affine actions. Abels, Margulis, and Soifer also contributed to related topics such as the Auslander conjecture (see Abels [2] for a survey). Danciger and Zhang [13] investigated affine deformations of Hitchin representations, showing that they do not exist. Danciger, Kassel, and Guéritaud [12] provided first examples of fundamental groups of closed manifolds acting properly and freely on affine spaces.
@@ 1.4. Outline
In Section 2, we will introduce affine structures and some basic facts on $\delta$-hyperbolic metric spaces. Suppose that $\Gamma_{M}$ is word-hyperbolic. Then $\hat{M}$ has a Gromov hyperbolic Riemannian metric by the Svarc-Milnor theorem. (see [35] and [45]. Svarc was a translated name of Albert Schwarz who is currently working as a professor in the University of California, Davis. He published this theorem earlier than Milnor in Russian.) We discuss the ideal boundary of $\hat{M}$, identifiable with the Gromov boundary, and the complete isometric geodesics in $\hat{M}$. We relate the space of complete isometric geodesics to the Gromov flow space.
@@
Section 4 has the main arguments. In Section 4.1, we will relate the partial hyperbolic property in the singular-value sense to that in the bundle sense. In Section 4.2, we show that the linear part of the holonomy representation of a complete affine manifold is $P$-Anosov if and only if it is partially hyperbolic. First, we do this when the linear holonomy group has a reductive Zariski closure. This involves the fact that $1$ has to be an eigenvalue of each element of the linear holonomy groups of complete affine manifolds. We relate it to singular values using the clever ideas of Danciger and Stecker, the spectrum theory of Breulliard-Sert [7], Benoist [3], [4], and the work of Potrie-Kassel [30]. We could generalize to the cases of nonreductive linear holonomy groups using the stability of the Anosov and partial hyperbolic properties. This will prove Theorem 1.1.

### v2

@@ Abstract.
Let $N$ be a manifold of dimension $m$ with a flat vector bundle given by a representation $\rho:\pi_{1}(N)\rightarrow{\mathsf{GL}}(n,{\mathbb{R}})$ where $\pi_{1}(N)$ is finitely generated. The holonomy group $\rho$ is a $k$-partially hyperbolic holonomy representation if the flat bundle pulled back over the unit tangent bundle of a sufficiently large compact submanifold of $N$ splits into expanding, neutral, and contracting subbundles along the geodesic flow, where the expanding and contracting subbundles are $k$-dimensional with $k<n/2$. Suppose that each element of $\rho(\pi_{1}(N))$ has an eigenvalue of norm $1$, or, alternatively, $\rho$ has some singular values of subexponential growth in terms of word length. We show that $\rho$ is a $P$-Anosov representation for a parabolic subgroup $P$ of ${\mathsf{GL}}(n,{\mathbb{R}})$ if and only if $\rho$ is a partially hyperbolic representation. We will primarily employ techniques from representation theory. As an application, we will show that the equivalence holds when $N$ is a complete affine $n$-manifold and $\rho$ is a linear part of the holonomy representation. This had never been done over the full general linear group.
@@ 1.1. Notation
Let $N$ be an $m$-manifold with a finitely generated fundamental group $\pi_{1}(N)$. Each flat vector bundle over $N$ can be described as ${\mathbb{R}}^{n}_{\rho}(N):=\tilde{N}\times{\mathbb{R}}^{n}/\sim$ where $(x,\vec{v})\sim(\gamma(x),\rho(\gamma)(\vec{v}))$ for $\gamma\in\pi_{1}(N)$ where $\rho:\pi_{1}(N)\rightarrow{\mathsf{GL}}(n,{\mathbb{R}})$ is a representation. A differential-geometry description of a flat bundle is a vector bundle with a flat linear connection.
Let $M$ be a compact submanifold (with or without boundary) in $N$. We assume that the inclusion map $M\hookrightarrow N$ induces a surjection on fundamental groups. The inverse image $\hat{M}$ of $M$ in $\tilde{N}$ is connected, and the deck transformation group $\Gamma_{M}$ of $\hat{M}$ is isomorphic to $\pi_{1}(N)$. In these cases, we refer to $M$, the map $\hookrightarrow N, M^\hat{M}$, the projection $p_{\hat{M}}:\hat{M}\rightarrow M$, and $\Gamma_{M}$ as a fundamental group surjective (FS) submanifold, an FS map, an FS cover, an FS covering map, and an FS deck transformation group, respectively.
@@
We will assume that $\partial M$ is convex and smooth in $N$. Let $d_{M}$ denote the path metric on $M$ induced from a Riemannian metric on $N$, and let $d_{\hat{M}}$ denote the path metric on $\hat{M}$ induced from it.
  • Let $\mathbf{U}M$ denote the space of direction vectors on $M$, i.e., the space of elements of the form $(x,\vec{v})$ for $x\in M$ and a unit vector $\vec{v}\in T_{x}M\setminus\{O\}$ with the projection $\pi_{\mathbf{U}M}:\mathbf{U}M\rightarrow M$.
@@
A complete isometric geodesic in $\hat{M}$ is a geodesic that is an isometry of $\mathbb{R}$} into $\hat{M}$ equipped with $d_{\hat{M}}$. Some but not all Riemannian geodesics are isometries. Actually, in geodesic metric spaces, these are identical to the notion of geodesics. Since we are working on Riemannian metric spaces, we use the adjectives. A complete isometric geodesic in $M$ is a geodesic that lifts to a complete isometric geodesic in $\hat{M}$. An isometric ray in $\hat{M}$ is an isometric geodesic defined on $0,\infty)$.
@@
$\gamma\cdot(x,\vec{v})=(\gamma(x),\rho(\gamma)(\vec{v}))\text{ for }x\in\mathbf{UC}\hat{M},\vec{v}\in{\mathbb{R}}^{n}.$
@@ 1.2. Main result
In this paper, we establish connections between the partial hyperbolicity in the context of dynamical systems and the Anosov properties.
@@
    ${\mathbb{V}}_{+}(x),{\mathbb{V}}_{0}(x),{\mathbb{V}}_{-}(x)\subset{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\mathbb{R}}^{n}_{\rho}(x)}$
    for the fiber ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\mathbb{R}}^{n}_{\rho}}(x)$ of ${\mathbb{R}}^{n}_{\rho}$ over each point $x$ of $\mathbf{UC}M$ invariant under the flow $\Phi_{t}$ of ${\mathbb{R}}^{n}_{\rho}$ over the geodesic flow $\phi_{t}$ of $\mathbf{UC}M$.
@@
    ${\mathbb{V}}_{+}(x),{\mathbb{V}}_{0}(x)$ and ${\mathbb{V}}_{-}(x)$ for each $x\in\mathbf{UC}M$ are independent subspaces, and their sum equals ${\mathbb{R}}^{n}_{\rho}(x)$.
@@
We additionally require that $\dim{\mathbb{V}}_{+}=\dim{\mathbb{V}}_{-}\geq 1$. Here $\dim{\mathbb{V}}_{+}$ is the partial hyperbolicity index of $\rho$ for the associated bundle decomposition ${\mathbb{V}}_{+},{\mathbb{V}}_{0},{\mathbb{V}}_{-}$ and $\dim{\mathbb{V}}_{+}=\dim{\mathbb{V}}_{-}<n/2$, so that $\dim{\mathbb{V}}_{0}\geq 1$. Furthermore, ${\mathbb{V}}_{0}$ is said to be the neutral subbundle of ${\mathbb{R}}^{n}_{\rho}$. (See Definition 1.5 of [10] or [6].) We use the words “in a bundle sense” only as a convenience of this paper.)
The definition depends on the Riemannian metric, but not on $\left|\mkern-2.0mu\left|\cdot\right|\mkern-2.0mu\right|_{{\mathbb{R}}^{n}_{\rho}}$. If $G_{M}$ is word-hyperbolic, then we can replace $\mathbf{UC}\hat{M}$ with a Gromov flow space which depends only on $G_{M}$ up to quasi-isometry. (See Theorem 2.6 and Gromov [23] or Mineyev [37].) Also, we leave as a question if Definition 1.1 forces $G_{M}$ to be word-hyperbolic.
@@
  • the linear part $\rho:\pi_{1}(N)=\Gamma_{M}\rightarrow{\mathsf{GL}}(n,{\mathbb{R}})$ is a partially hyperbolic representation in a bundle sense over $M$.
We denote by $a_{1}(A),\dots,a_{n}(A)$ the nonincreasing set of singular values of an $n\times n$-matrix $A$. We denote by $\lambda_{1}(A),\dots,\lambda_{n}(A)$ the nonincreasing set of norms of eigenvalues of an $n\times n$-matrix $A$. (Here, the multiplicity is allowed.)
@@
$\theta=\left\{\log\lambda_{i_{1}}-\log\lambda_{i_{1}+1},\dots,\log\lambda_{i_{m}}-\log\lambda_{i_{m}+1}\right\},1\leq i_{1}<\cdots<i_{m}\leq n-1,$
of ${\mathsf{GL}}(n,{\mathbb{R}})$, the parabolic subgroup $P_{\theta}$ (resp. $\left.P_{\theta}^{-}\right)$ is the subgroup of block upper (resp. lower) triangular matrices in ${\mathsf{GL}}(n,{\mathbb{R}})$ with square diagonal blocks of sizes $i_{1},i_{2}-i_{1},\dots,i_{m}-i_{m-1},n-i_{m}$ respectively. We denote by $\theta^{*}$ the image of $\theta$ under the map $i\mapsto n-i$ for $i=1,\dots,n-1$.
For a finitely generated group $G$, we denote by $w(g)$ the word length of $g$ in terms of a fixed finite generating set $S$ in $G$.
@@
Let $N$ be an $m$-manifold with a finitely generated word-hyperbolic fundamental group. Let $\rho:\pi_{1}(N)\rightarrow{\mathsf{GL}}(n,{\mathbb{R}})$ be a representation. Assume one of the following:
  • (i)
    $|\lambda_{i_{g}}(\rho(g))|=1$ for all $g\in\pi_{1}(N)$ where the index $i_{g}\in\{1,\dots,n\}$ may depend on $g$.
  • (ii)
    For any sequence $g_{i}\in\pi_{1}(N)$, if $w(g_{i})\rightarrow\infty$, then $\frac{|\log a_{j_{g_{i}}}(\rho(g_{i}))|}{w(g_{i})}\rightarrow 0$ for some choices of the indices $j_{g_{i}}\in\{1,\dots,n\}$ depending on $g_{i}$.
Then $\rho$ is $P$-Anosov for any parabolic subgroup $P$ of ${\mathsf{GL}}(n,{\mathbb{R}})$ with $P=P_{\theta}$ for $\theta=\theta^{*}$ containing $\log\lambda_{k}-\log\lambda_{k+1},k\leq n/2$ if and only if $\rho$ is partially hyperbolic of the index $k$ in a bundle sense. In these cases $k<n/2$.
We say that $\rho$ has subexponential growth for some singular values if (ii) holds. By Lemma 2.11 of [7] applied to the standard representation of ${\mathsf{GL}}(n,{\mathbb{R}})$, (ii) implies (i); however, we wrote out (ii) for the purpose of the usefulness of the result. Also, the convergence in (ii) is uniform with respect to the word length since we can find a seqeunce not converging to 0 otherwise.
It is easy to see that the conclusion is true for representations to ${\mathsf{SL}}_{\pm}(n,{\mathbb{R}})$ and for $k$ which is the least element of $\theta$ without our conditions. If $\pi_{1}(N)$ is not simple, we can multiply any representation by a representation to $\mathbb{R}$$^{+}$ so that a partial hyperbolicity cannot hold. We can ask if an Anosov representation is always partially hyperbolic up to multiplying by a positive scalar representation to $\mathbb{R}$}. We cannot answer this question since controlling the singular values of representation elements is very hard in general, and the only related results that we know of are due to Abel-Margulis-Soifer [1], which is not applicable here as far as we know. See Kassel-Potrie [30] for some discussions. However, only cases that are left are for representations whose singular values behave exponentially.
@@
Suppose that $N$ is a complete affine $n$-manifold with a finitely generated word-hyperbolic fundamental group. Let $\rho:\pi_{1}(N)\rightarrow{\mathsf{GL}}(n,{\mathbb{R}})$ be a linear part of the affine holonomy group. Then the conclusions of Theorem 1.1 hold.
@@
We wish to approach the conjecture that a closed complete affine manifold cannot have a word-hyperbolic fundamental group. One approach considered is using Oseledets-type theorems (See Chapter 8 of [32]). However, generalizing our results to measurable settings is currently hindered by numerous technical issues. Discussions on these matters took place with researchers such as Nicholas Tholozan at BIRS, Banff, during the winter of 2020.
@@ 1.3. Some historical comments on complete affine manifolds
Further developments include Ghosh [15, 17, 16, 18], Ghosh-Treib [19], and Smilga [39, 40, 41, 42, 43], who generalized these results and produced higher-dimensional examples of proper and free affine actions. Abels, Margulis, and Soifer also contributed to related topics such as the Auslander conjecture (see Abels [2] for a survey). Danciger and Zhang [13] investigated affine deformations of Hitchin representations, showing that they do not exist. Danciger, Kassel, and Guéritaud [12] provided the first examples of fundamental groups of closed manifolds acting properly and freely on affine spaces.
@@ 1.4. Outline
In Section 2, we will introduce affine structures and some basic facts on $\delta$-hyperbolic metric spaces. Suppose that $\Gamma_{M}$ is word-hyperbolic. Then $\hat{M}$ has a Gromov hyperbolic Riemannian metric by the Svarc-Milnor theorem. (See [36] and [46]. Švarc is a transliteration of Albert Schwarz’s name. He published this theorem in Russian prior to Milnor. Currently, he is a professor at the University of California, Davis.) We discuss the ideal boundary of $\hat{M}$, identifiable with the Gromov boundary, and the complete isometric geodesics in $\hat{M}$. We relate the space of complete isometric geodesics to the Gromov flow space.
@@
Section 4 has the main arguments. In Section 4.1, we will relate the partial hyperbolic property in the singular-value sense to that in the bundle sense. In Section 4.2, we show that the linear part of the holonomy representation of a complete affine manifold is $P$-Anosov if and only if it is partially hyperbolic. First, we do this when the linear holonomy group has a reductive Zariski closure. We use the fact that $1$ has to be an eigenvalue of each element or the subexponential growth of some singular values of the linear holonomy groups. We relate it to singular values using the clever ideas of Danciger and Stecker, the spectrum theory of Breuillard-Sert [7], Benoist [3], [4], and the work of Potrie-Kassel [31]. We could generalize to the cases of nonreductive linear holonomy groups using the stability of the Anosov and partial hyperbolic properties. This will prove Theorem 1.1.
In Section 5, we apply these results to complete affine manifolds. proving Corollary 1.2.